An urn contains 5 black balls and 2 white balls. Four balls are removed at random from the urn. The random variable X is the number of white balls among the four balls that have been removed. Find P(X = 2), i.e. the probability that X = 2.

A. 1/7

B. 2/7

C. 4/21

D. 1/42

E. None of the above answers is correct

Having a really hard time remembering how to do these types of problems! Please be detailed in how the answer is found. THANK YOU! :D
Answer #1

We have 5 black balls and 2 white balls

Total 7 balls in urn

We have to take 4 balls randomly out of 7

I.e. 7c4

Let X is number of white balls

Find P(X=2)

We want to take 2 white balls and 2 blacks balls

(2c2 * 5c2)/ (7c4) = 10/35 = 2/7

Answer :- [ B ]